Circular No. 6625

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) D. Mehringer, M. Gardner, and D. Benford, Caltech Submillimeter Ob-

servatory (CSO); D. Bockelée-Morvan, Observatoire de Paris (OP); D. C. Lis and T. G. Phillips, CSO; D. Despois, Observatoire de Bordeaux; and N. Biver, H. Rauer, P. Colom, J. Crovisier, and D. Gautier, OP, report the first groundbased spectral identification of H_{3}O^{+} in a comet using the CSO: “On Apr. 6 and 10 UT, we detected the 3(2)–2(2) line of H_{3}O^{+} at 364.797 GHz in comet C/1995 O1. The integrated line area on the antenna temperature scale is 0.248 ± 0.046 K km s^{−1} (average of the two days). The line is approximately centered on the comet’s rest velocity, and its full width at half maximum is ∼ 3 km/s. The H_{3}O^{+} line at 307.192 GHz was not detected with a 3σ upper limit of 0.09 K km s^{−1}.”

M. J. Mumma and V. A. Krasnopolsky, Goddard Space Flight Center, NASA; M. Abbott and B. C. Flynn, Center for Extreme Ultraviolet Astrophysics, Berkeley; D. K. Yeomans, Jet Propulsion Laboratory; P. D. Feldman, Johns Hopkins University; and C. B. Cosmovici, Istituto di Fisica dello Spazio Interplanetario, report: “Soft x-rays, He I (58.4 nm), and O II (53.8 nm) were detected in comet C/1995 O1 during 1996 Sept. 14–19 UT (r = 3.07 AU) with the EUVE orbiting observatory. The intensity in x-rays (70–180 eV) was 8 × 10^{24} photons/s in an aperture of radius 400 000 km, and the central brightness was displaced from the nucleus by 140 000 ± 60 000 km in the sky plane. A region of extended x-ray emission is also seen, and this is anticorrelated with the dust jets imaged simultaneously in the visible. He I 58.4-nm and O II 53.8-nm lines were detected with S/N = 4. The absence of Ne lines shows a depletion of Ne in the comet by more than 25 times relative to the solar abundance.”

Yu. N. Gnedin and T. M. Natsvlishvili, Central Astronomical Observatory, Pulkovo; and V. D. Bychkov and V. P. Romanenko, Special Astrophysical Observatory, report preliminary results of polarimetric observations of this comet with a 1-m telescope (+ MINIPOL polarimeter and 6′′ diaphragm) during Mar. 28–Apr. 1. The amount of polarization is similar in the coma, the dust tail, and the ion tail, varying generally from 5–10 percent in U, 10–19 in B, and 7–17 in both R and I. The V polarization is generally < 1 percent in the coma, increasing to 10–20 percent in both the dust and ion tails. The high degree of polarization and associated strong wavelength dependence (especially at the center of the comet on Mar. 28 and 29) give strong evidence for molecular-band polarization. The existence of CN, CO^{+}, and H_{2}O^{+} species are considered most probable.

1997 April 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green